Determine the equations of the tangent and normal to:



1. π‘‡β„Žπ‘’ π‘π‘Žπ‘Ÿπ‘Žπ‘π‘œπ‘™π‘Ž 𝑦² = 6π‘₯ βˆ’ 3 π‘π‘’π‘Ÿπ‘π‘’π‘›π‘‘π‘–π‘π‘’π‘™π‘Žπ‘Ÿ 𝑙𝑖𝑛𝑒 π‘₯ + 3𝑦 = 7



2. π‘‡β„Žπ‘’ 𝑒𝑙𝑙𝑖𝑝𝑠𝑒 π‘₯Β² + 2𝑦² = 9 π‘π‘’π‘Ÿπ‘π‘’π‘›π‘‘π‘–π‘π‘’π‘™π‘Žπ‘Ÿ π‘‘π‘œ π‘‘β„Žπ‘’ 𝑙𝑖𝑛𝑒 4π‘₯ βˆ’ 𝑦 = 6



3. π‘‡β„Žπ‘’ π‘π‘’π‘Ÿπ‘£π‘’ (π‘₯ βˆ’ 𝑦) Β² + 4π‘₯ βˆ’ 1 = 0 π‘Žπ‘‘ (1,0)



4. π‘‡β„Žπ‘’ π‘π‘’π‘Ÿπ‘£π‘’ 𝑦 = 2π‘₯/π‘₯+2 π‘Žπ‘‘ π‘₯ = 2




Locate the critical points and determine whether the given functions are either



maxima or minima using First Derivate Test and Second Derivative Test:



5. 𝑦 = π‘₯ Β²(π‘₯ βˆ’ 2)Β²



6. 𝑦 = π‘₯(π‘₯Β² βˆ’ 6π‘₯ βˆ’ 15) + 50



7. 9aΒ³y = x(4a βˆ’ x)Β³

1
Expert's answer
2022-01-24T03:11:15-0500

"1)\qquad\\
\text{We are given,}\\
\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad y^2=6x-3\\
\text{From the given line } x+3y=7\\
\text{We have}\\
\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad y=-\frac{x}{3}+\frac{7}{3}\\
\text{Therefore, the Slope of this line is }m=-\frac{1}{3}.\\
\qquad \text{The Slope perpendicular to the Slope of the given line (m$\perp$) is }\mathbf{m_{\perp}}=3.\\
\text{Next, differentiating }y^2=6x+3 \text{ implicitly wrt } x:\\
\Rightarrow2yy\prime=6\\
\Rightarrow y\prime=\frac{3}{y}=\mathbf{m_{tan}}, \text{where $\mathbf{m_{tan}}$ is the Slope of a tangent line to the curve } y^2=6x-3.\\
\text{Next, } \mathbf{m_{\perp}}=\mathbf{m_{tan}}\\
\Rightarrow3=\frac{3}{y}\\
\Rightarrow y=1\\
\text{Therefore }y^2=6x-3\\
\Rightarrow 1^2=6x-3\Rightarrow x=\frac{2}{3}\\
\text{Therefore, the eqn. of the tangent passing through ($\frac{2}{3},1$) is}\\
\frac{y-1}{x-\frac{2}{3}}=m_\perp=3\\
\Rightarrow y=3x-1\\
\text{Therefore, the eqn. of the tangent line is}\\
\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad y=3x-1\\
\qquad\\
\text{Next,}\\
\mathbf{m_{normal}}=-\frac{1}{\mathbf{m_{tan}}}=-\frac{y}{3}\\
\mathbf{m_{normal}}=\mathbf{m_{\perp}}\\
\Rightarrow3=-\frac{y}{3}\Rightarrow y=-9\\
\text{So, }y^2=6x-3\Rightarrow(-9)^2=6x-3\Rightarrow x=14\\
\text{Therefore, the eqn. of the normal passing through (14,-9) is}\\
\frac{y+9}{x-14}=m_\perp=3\Rightarrow y=3x-51.\\
\quad\\
\quad\\
\quad\\
2)\qquad\\\text{We are given,}\\
\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad x^2+2y^2=9\\
\text{From the given line } 4x-y=6\\
\text{We have}\\
\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad y=4x-6\\
\text{Therefore, the Slope of this line is }m=4.\\
\qquad \text{The Slope perpendicular to the Slope of the given line (m$\perp$) is }\mathbf{m_{\perp}}=-\frac{1}{4}.\\
\text{Next, differentiating }x^2+2y^2=9 \text{ implicitly wrt } x:\\
\Rightarrow 2x+4yy\prime=0\\
\Rightarrow y\prime=\frac{-x}{2y}=\mathbf{m_{tan}}, \text{where $\mathbf{m_{tan}}$ is the Slope of a tangent line to the curve } x^2+2y^2=9.\\
\text{Next, } \mathbf{m_{\perp}}=\mathbf{m_{tan}}\\
\Rightarrow -\frac{1}{4}=-\frac{x}{2y}\\
\Rightarrow y=2x\\
\text{Therefore }x^2+2y^2=9\\
\Rightarrow x^2+2(2x)^2=9\Rightarrow x=-1,1\\
\text{At x=-1, y=-2}\\
\text{At x=1, y=2}\\
\quad\text{The eqn. of the tangent passing through ($-1,-2$), is}\\
\frac{y+2}{x+1}=m_\perp=-\frac{1}{4}\\
\Rightarrow y=-\frac{x}{4}-\frac{9}{4}\\
\quad\text{Also, the eqn. of the tangent passing through ($1,2$), is}\\
\frac{y-2}{x-1}=m_\perp=-\frac{1}{4}\\
\Rightarrow y=-\frac{x}{4}+\frac{9}{4}\\
\text{Therefore, the eqn. of the tangent lines are}\\
\qquad\qquad\qquad\qquad\qquad\qquad y=-\frac{x}{4}-\frac{9}{4},\ \ y=-\frac{x}{4}+\frac{9}{4}\\
\qquad\\
\text{Next,}\\
\mathbf{m_{normal}}=-\frac{1}{\mathbf{m_{tan}}}=\frac{2y}{x}\\
\mathbf{m_{normal}}=\mathbf{m_{\perp}}\\
\Rightarrow -\frac{1}{4}=\frac{2y}{x}\Rightarrow y=-\frac{x}{8}\\
\text{So, }x^2+2y^2=9\Rightarrow x^2+2(-\frac{x}{8})^2=9\Rightarrow x=\pm\sqrt{\frac{96}{11}}\\
\text{At }x=\sqrt{\frac{96}{11}}, y=-\sqrt{\frac{3}{22}}\\
\text{At }x=-\sqrt{\frac{96}{11}}, y=\sqrt{\frac{3}{22}}\\
\quad\text{The eqn. of the tangent passing through ($\sqrt{\frac{96}{11}},-\sqrt{\frac{3}{22}}$), is}\\
\frac{y+\sqrt{\frac{3}{22}}}{x-\sqrt{\frac{96}{11}}}=m_\perp=-\frac{1}{4}\\
\Rightarrow y=-\frac{x}{4}-\sqrt{\frac{6}{11}}\\
\quad\text{Also, the eqn. of the tangent passing through ($-\sqrt{\frac{96}{11}},\sqrt{\frac{3}{22}}$), is}\\
\frac{y-\sqrt{\frac{3}{22}}}{x+\sqrt{\frac{96}{11}}}=m_\perp=-\frac{1}{4}\\
\Rightarrow y=-\frac{x}{4}+\sqrt{\frac{6}{11}}\\
\qquad\qquad\qquad\qquad\qquad\qquad y=-\frac{x}{4}-\sqrt{\frac{6}{11}},\ \ y=-\frac{x}{4}+\sqrt{\frac{6}{11}}\\
\quad\\
\quad\\
\quad\\
3)\quad\\
\text{We are given }(x-y)^2+4x-1=0\\
\text{differentiating }(x-y)^2+4x-1=0 \text{ implicitly wrt x;}\\
\Rightarrow2(x-y)(-y\prime)+4=0\\
\Rightarrow y\prime=\frac{2}{x-y}\\
\text{At } (1,0);\\
m_{tan}=y\prime=\frac{2}{1-0}=2\\
\text{Thus, the equation of the tangent of the curve at (1, 0) is;}\\
\frac{y-0}{x-1}=2=m_{tan}\Rightarrow y=2(x-1)\\
\quad\\
\text{Next, }m_{tan}\times m_{normal}=-1\Rightarrow m_{normal}=\frac{-1}{2}\\
\text{Thus, the equation of the normal of the curve at (1, 0) is;}\\
\frac{y-0}{x-1}=\frac{-1}{2}=m_{normal}\Rightarrow y=-\frac{x}{2}+\frac{1}{2}\\
\quad\\
\quad\\
\quad\\
4)\quad\\
\text{We are given }y=\frac{2x}{x+2}\\
\text{differentiating }y=\frac{2x}{x+2} \text{ implicitly wrt x;}\\
\Rightarrow y\prime=\frac{4}{(x+2)^2}\\
\text{At }x=2;\\
m_{tan}=y\prime=\frac{1}{4}, y=1\\
\text{Thus, the equation of the tangent of the curve passing through (2, 1) is;}\\
\frac{y-1}{x-2}=\frac{1}{4}=m_{tan}\Rightarrow y=\frac{x}{4}+\frac{1}{2}\\
\quad\\
\text{Next, }m_{tan}\times m_{normal}=-1\Rightarrow m_{normal}=-4\\
\text{Thus, the equation of the normal of the curve at (2, 1) is;}\\
\frac{y-1}{x-2}=-4=m_{normal}\Rightarrow y=-4x+9\\
\quad\\
\quad\\
\quad\\
5)\\
y=x^2(x-2)^2, \text{differentiating yields:}\\y\prime=4x(x-1)(x-2)\\
\quad\text{At critical point, }y\prime=0 \Rightarrow x=0,1,2 \text{ are the critical points.}\\
\text{\textbf{FIRST DERIVATIVE TEST:}}\\
y\prime(-1)=-24\Rightarrow decreasing\ on\ (-\infty,0) \\ 
y\prime(\frac{1}{2})=\frac{3}{2}\Rightarrow increasing\ on\ (1,2)\\
y\prime(\frac{3}{2})=-\frac{3}{2}\Rightarrow decreasing\ on\ (0,1) \\
y\prime(3)=24\Rightarrow increasing\ on\ (2,\infty)\\
\text{Since switching from decreasing to increasing and increasing to decreasing at a point}\\
\Rightarrow \text{the point is a local minimum point and local maximum point respectively, thus}\\
\text{x = 1 is a local maximum point and x = 0, 2 are local minimum points.}\\
\text{Thus, the function has local maximum at (1, 1) and local minimum at (2, 0),(0, 0)}.\\
\quad\\
\text{\textbf{SECOND DERIVATIVE TEST:}}\\
y\prime\prime=12x^2-24x+8\\
\text{Plug in the critical numbers:}\\
y\prime\prime(0)=8\Rightarrow0 \text{ is a minimum point}\\
y\prime\prime(1)=-4\Rightarrow1 \text{ is a maximum point}\\
y\prime\prime(2)=8\Rightarrow0 \text{ is a minimum point}\\
\text{Thus, the function has local maximum at (1, 1) and local minimum at (2, 0),(0, 0)}.\\
\quad\\
\quad\\
\quad\\
6)\\
y=x^3-6^2-15x+50, \text{differentiating yields:}\\y\prime=3(x-5)(x+1)\\
\quad\text{At critical point, }y\prime=0 \Rightarrow x=5,-1 \text{ are the critical points.}\\
\text{\textbf{FIRST DERIVATIVE TEST:}}\\
y\prime(-2)=21\Rightarrow increasing\ on\ (-\infty,-1) \\ 
y\prime(0)=-15\Rightarrow decreasing\ on\ (-1,5)\\
y\prime(9)=120\Rightarrow increasing\ on\ (5,\infty) \\
\text{Since switching from decreasing to increasing and increasing to decreasing at a point}\\
\Rightarrow \text{the point is a local minimum point and local maximum point respectively, thus}\\
\text{x = -1 is a local maximum point and x = 5 is a local minimum point.}\\
\text{Thus, the function has local maximum at (-1, 58) and local minimum at (5, -50)}.\\
\quad\\
\text{\textbf{SECOND DERIVATIVE TEST:}}\\
y\prime\prime=6x-12\\
\text{Plug in the critical numbers:}\\
y\prime\prime(-1)=-18\Rightarrow0 \text{ is a maximum point}\\
y\prime\prime(5)=18\Rightarrow1 \text{ is a minimum point}\\
\text{Thus, the function has local maximum at (-1, 58) and local minimum at (5, -50)}.\\
\quad\\
\quad\\
\quad\\
7)\\
y=\frac{x(4a-x)^3}{9a^3}, \text{differentiating yields:}\\y\prime=\frac{(4a-x)^2(4a-4x)}{9a^3}\\
\quad\text{At critical point, }y\prime=0 \Rightarrow x=a, 4a \text{ are the critical points.}\\
\quad\\
\text{Determine the 2nd derivative:}\\
9a^3 y'' = -12(4a - x)(2a - x)\\
\quad\\
x = 4a\Rightarrow\\
y=0\\
\text{Check if Maxima or Minima.}\\
9a^3 y''(4a) = -12(4a - 4a)(2a - 4a)\Rightarrow y'' = 0 \ (\text{inflection point})\\
\text{Hence, (4a, 0) is neither maximum nor minimum}\\
\quad\\
x=a \Rightarrow y=3a\\
\text{Check if Maxima or Minima}\\
y''(a) = (-)\ (\text{the curve is concave downward})\\
\text{Hence, (a, 0) is maximum.}"